OpenStudy (anonymous):

4. Solve the proportion using number sense. 2/8 X/48 A) 6 B) 12 C) 16 D) 8 5. Solve the proportion using cross products. 18/20 K/110 A) 99 B) 122 C) 3.3 D) 2.9 Yeah im really asking this xD im tired ok..........................

OpenStudy (anonymous):

2/8= x/48 crossmultiply 8x = 96 x= 96/8 = ?

OpenStudy (anonymous):

and use that method for the next question also! :-)

OpenStudy (anonymous):

;-; i know how to cross multiply!! ;-; Im going to the ER though, n i want my math work done........... before i get there ;-;... @mangorox

OpenStudy (anonymous):

for the first question 96/8 = 12 < answer

OpenStudy (anonymous):

For the second question: 18/20 = k/110 20k = 1980 x = 1980/20 = 99<answer

OpenStudy (anonymous):

Thanks so much... Sorry i would be doing this but its kinda hard to think right now ):

OpenStudy (anonymous):

That's ok!

OpenStudy (anonymous):

c: